OpenStudy (anonymous):

When Sophie and her dog go for a walk, Sophie keeps track of the time and distance they have walked. They always walk at the same constant rate so that the distance they go is proportional to the amount of time they have been walking. If it takes 20 minutes to walk around the park four times, how many minutes will it take them to walk around the park two times?

OpenStudy (texaschic101):

20/4 = x/2 (20 minutes at 4 times = x minutes at 2 times) 4x = (2)(20) result of cross multiplying) 4x = 40 x = 10 ANSWER : It would take 10 minutes to walk around the park 2 times.